OpenStudy (anonymous):

How would I graph y=2?

OpenStudy (amistre64):

draw a stright line thru y = 2

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

This may be a stupid question, but what would the other coordinate be?

OpenStudy (amistre64):

there is no other coordinate; and value of x will work. (x,2) is the solution. and that is a straight line from left to right thru the y = 2 mark on the y axis

OpenStudy (anonymous):

Thank you again!

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

Sorry to keep this up, it is telling me to graph the equation y=2, and it is making me put in two sets of coordinates, I'm really confused sorry

OpenStudy (amistre64):

use any number for x and keep y at 2 (5,2) (-5,2) (0,2) (-26/3 , 2) (.56,2) (any#,2)

OpenStudy (anonymous):

Ok that makes sense, thank you again

OpenStudy (amistre64):

y = 0x +2 for any value of 'x' ; what does y=?

OpenStudy (anonymous):

2?

OpenStudy (amistre64):

exactly; and we can prove that by plugging in any value for 'x and we get a 2 as a result right?

OpenStudy (amistre64):

y = 0(34) + 2 = 2 y = 0(13764.2846) + 2 = 2 :)

OpenStudy (anonymous):

Yes cause anything multiplied by 0 is 0 and that just leaves the two?

OpenStudy (amistre64):

correct :)

OpenStudy (anonymous):

Ok You are my hero thank you so much

OpenStudy (amistre64):

you are welcome :)